### v1

# Lifetime-Limited Linewidth Measurements of the 3C and 3D Soft X-ray Transitions in Ni XIX

Abstract
We used the monochromatic soft X-ray beamline P04 at the synchrotron-radiation facility PETRA III to resonantly excite the strongest transitions in neon-like Ni xix ions, and , respectively dubbed 3C and 3D, achieving a resolving power of 15,000 and signal-to-background ratio of 30. We obtain their natural linewidths, with an accuracy of better than 10%, as well as the oscillator-strength ratio = 2.51(11) from analysis of the resonant fluorescence spectra. These results agree with those of previous experiments, earlier predictions, and our own advanced calculations.
I Introduction
Because of their closed-shell ground-state configuration, neon-like ions are prevalent in plasmas across a broad range of temperatures. Their strong spectral lines provide a wealth of diagnostic capabilities, including temperature, density, optical depth, and ultraviolet field intensity. Their spectra were extensively studied in investigations of the Sun and other celestial bodies with the Chandra and XMM-Newton observatories Parkinson (1973); Smith et al. (1985); Schmelz et al. (1992); Waljeski et al. (1994); Phillips et al. (1996); Mauche et al. (2001); Behar et al. (2001); Doron and Behar (2002); Xu et al. (2002a); Gu (2003); Paerels and Kahn (2003); Pradhan and Nahar (2011); Beiersdorfer et al. (2018); Gu et al. (2019, 2020); Grell et al. (2021). The spectrum of Fe XVII is the most commonly studied in the soft x-ray band due to the high cosmic abundance of iron and the easily accessible energy of its transitions, while Ni XIX is the second-most abundant such species.
Despite the apparent simplicity of their electronic structure, systematic discrepancies in the intensities of their strongest emission lines were long noted between theory, astrophysical observations, and laboratory measurements (see Brown (2008) and references therein). Recently, we solved in Ne-like Fe XVII a problem that persisted for several decades: the oscillator-strength ratio between its resonance (3C: ) and intercombination (3D: ) lines consistently deviated from predictions Brown (2008). Previous experiments on this topic had suffered from both known and unexpected systematic uncertainties Brown et al. (1998, 2001a, 2001b, 2006); Beiersdorfer et al. (2002); Gillaspy et al. (2011); Beiersdorfer et al. (2004, 2017); Shah et al. (2019); Bernitt et al. (2012); Oreshkina et al. (2014); Loch et al. (2015); Kühn et al. (2020). Our resonant excitation scheme using synchrotron radiation with much increased resolution and signal-to-noise ratio finally brought experiment and theory into agreement Kühn et al. (2022).
In most measurements of Fe XVII, an inner-shell satellite from Na-like ions (line C: ) blended with the Ne-like line 3D, leading to potential systematic errors in the line ratio Brown et al. (2001b). Experiments in which 3C and 3D were excited by electron impact, and which had sufficient spectral resolving power to detect other lines of Na-like Fe, allowed correction of the strength of line 3D for contamination from line C, as well as optimization of the experimental conditions to minimize production of that undesired charge state. Photoexcitation measurements suffer from a second, subtler effect: the strong autoionizing branch of the upper level of line C continuously feeds the Ne-like ground state. In these experiments with insufficient resolution to split lines C and 3D, the resulting population transfer severely affected the apparent 3C/3D line-intensity ratio Wu and Gao (2019). Only the most recent photoexcitation measurements mentioned above Kühn et al. (2022), with a resolving power of 20,000, could reduce this detrimental transfer effects, reducing the effective overlap between lines C and 3D to less than 1%.
Another approach to understanding issues in the 3C/3D ratio in Ne-like iron is to study the same ratio along the isoelectronic sequence as a function of atomic number . This has two advantages: first, any systematic errors peculiar to a single experiment are revealed as outliers; second, the scaling with atomic number of any deviation from predictions can guide future theoretical investigations. 3C/3D ratios for Ne-like ions ranging from Cr XV to Kr XXVII () were measured at the Lawrence Livermore National Laboratory (LLNL) using an electron beam ion trap (EBIT) equipped with crystal spectrometers Brown et al. (2001c), showing systematic departures from theory from 10–20%. Extending our photoexcitation experiment to Ne-like Ni, where the lines 3D and C are much further apart than in Fe XVII, would fully suppress the undesired 3D-C overlap. This, the closeness of nickel to iron in , and the astrophysical importance of nickel motivated our present measurement in Ne-like Ni XIX ions. Because of the lower chemical abundance of nickel, its L-shell lines are weaker than those of iron in astrophysical sources. Ni XIX is nevertheless extremely useful for understanding the spectra of the solar and stellar coronae Loulergue and Nussbaumer (1975); Hutcheon et al. (1976) and for determinations of abundances, as well as plasma temperature and density Gu et al. (2004). In many high-energy-density plasmas Rogers and Iglesias (1994); Seaton et al. (1994); Beiersdorfer et al. (1996), 3C and 3D are stronger than the other L-shell transitions and affect the Rosseland mean opacity, for which recent studies at temperatures akin to stellar interiors disagreed by 10-20% from models for Ni Nagayama et al. (2019). Thus, an experimental benchmark is also needed for validating the underlying atomic data in opacity models, and could help clarifying the iron-opacity problem Bailey et al. (2007); Fontes et al. (2015).
Ability of atomic methods to accurately predict core parameters, such as transition energies, transition rates, and subsequently derived values, e.g., collisional cross sections, critically depends on the benchmark tests where predicted values are compared against the experiment. The most readily accessible properties are transition energies, where most accurate experimental data can be obtained. Comparing theoretical and experimental energies is an excellent start on testing the atomic methods, as discrepancies in the energies immediately point to the method deficiencies. However, the agreement of the energies does not predict a high level of accuracy of all the other properties. Energy comparisons do not account for different dependence of the atomic properties on the distance from the nucleus, additional correlation corrections specific to the transition operators, and subtler effects of the configuration mixing. Thus, to fully validate atomic structure calculations, experiments measuring natural linewidths and lifetimes are essential.
Lifetimes of excited HCI from the optical to the X-ray domain have been measured from decades using among others accelerators, storage rings, and electron beam ion traps (EBIT). At accelerators, beam-foil techniques cover a few ns to hundreds of fs range Betz et al. (1974); Träbert (2005, 2010), relying on spatially resolving X-ray emission following the passage of a fast ion through a thin foil. However, complexities arising from multiple excitations and charge-state distribution limits in general the accuracy of such data. Kingdon traps were also used for some lifetime measurements on ions Moehs et al. (2000, 2001); Smith et al. (2005). Storage rings have allowed for many accurate lifetime studies up to the range of seconds, e. g. in Refs. Mannervik et al. (1997); Träbert et al. (2002); Träbert (2002); Träbert et al. (2003, 2004, 2006, 2012, 2012a, 2012b); Träbert (2022, 2023) upon injection of excited HCI from external sources Habs et al. (1989); Doerfert et al. (1997), and in the optical range also using re-excitation of the circulating ions with lasers Klaft et al. (1994); Seelig et al. (1998); Lochmann et al. (2014). At EBITs, lifetimes ranging from ms down to ns are accurately measured by monitoring the decay of fluorescence after pulsed excitation, yielding uncertainties as low as 0.15% Crespo López-Urrutia et al. (1998); Beiersdorfer et al. (2003); Crespo López-Urrutia et al. (2006); Crespo López-Urrutia and Beiersdorfer (2010); Beiersdorfer et al. (2016). For the fs range, natural linewidths were accessed with high-resolution crystal spectrometers using standalone EBITs Beiersdorfer et al. (1996), or combined with synchrotron-radiation excitation Rudolph et al. (2013); Steinbrügge et al. (2015, 2022). Yet, modeling line profiles remains difficult, substantially limiting the achievable accuracy.
In this Letter, we present resonant photoexcitation of trapped Ni XIX ions at the monochromator beamline P04 at the PETRA III facility, with a focus on the two strong emission lines 3C and 3D. Several experimental improvements raised our resolving power to 15,000 and the signal-to-noise ratio to 30, enabling an accurate determination of the 3C/3D oscillator-strength ratio with approximate statistical and total uncertainties of 0.5% and 4.5%, respectively. Our observed line shapes yield the absolute natural linewidths, lifetimes, and oscillator strengths for 3C and 3D lines with an accuracy better than 10%. All measurements agree well with our new predictions.
II Measurements and Analysis
We carried out the experiment with PolarX-EBIT Micke et al. (2018), which is dedicated to the study of interactions of trapped highly charged ions (HCI) with photons from external sources, at the P04 beamline Viefhaus et al. (2013) of the PETRA III synchrotron-radiation facility. An off-axis electron gun emits a nearly mono-energetic electron beam that is compressed to a diameter of less than 100 m by an 870 mT magnetic field generated by an array of permanent magnets. Precursor atoms of Ni were brought in as a tenuous molecular beam of nickelocene (bis(cyclopentadienyl)nickel ) injected into the trap region through a two-stage differential pumping system. Successive electron impacts ionize Ni atoms to the charge state of choice. The ions are confined radially by the negative space-charge potential produced by the 5.5 mA and 1.1 keV electron beam, and axially by the 10 V potential difference given to the drift tubes before and after the central one. At the soft X-ray beamline P04, an APPLE II undulator generates circularly polarized photons which are monochromatized with a variable line-spacing platinum-coated 1200 l/mm grating Viefhaus et al. (2013). A Kirkpatrick-Baez (KB) mirror system refocuses the photon beam at the position of the PolarX trap region a few meters downstream. The focus diameter there is slighlty smaller than the ion cloud itself, which is approximately 200 micrometers broad. The photon beam enters PolarX from the side where the off-axis electron gun is mounted, and propagates along its longitudinal axis, defined by the magnetic field and the narrow electron beam that it guides and compresses. The photon beam is focused for maximum overlap onto the approximately 16 mm long cloud of Ni ions confined within the central trap electrode. Upon excitation, the resulting fluorescence is detected by two identical silicon drift detectors (SDD) with a resolving power of 10% at 1 keV mounted at right angles to the photon beam on the top and the side of the trapping region. After passing the trap region, the photon beam exits unimpeded PolarX through its collector, and enters a downstream beamline, where we measure its intensity.
For production of Ni XIX ions, the electron-beam energy must exceed 607 eV. This also leads to electron-impact-induced fluorescence from dielectronic recombination as well as resonant and direct excitation Shah et al. (2019) that is much stronger than the sought-after photoexcitation signal. We address this by periodically switching within a few microseconds the electron-beam energy between two optimized values, breeding for 200 ms Ni XIX ions at 1090 eV and detecting for 50 ms their fluorescence at 280 eV. This suppresses in-band background photons induced by electron-impact processes, resulting in a cleaner photoexcitation signal Kühn et al. (2022). The residual background results from electron impact production of few-hundred-eV photons which partly blend in the SDD with 3C and 3D due to its finite energy resolution and the width of the region of interest. Optimization of the ion-breeding duty cycle allowed us to further narrow the monochromator exit slit width to 25 m while keeping a strong fluorescence signal and achieving a robust signal-to-background ratio close to 30. The narrow slit allows the resolving power of the monochromator to reach values up to 15,000 for the Ni XIX 3C and 3D transitions.
We scanned a range of 500 meV around the known centroid energies Gu et al. (2004, 2007) in 100 steps of 10 meV each, while integrating the photoexcitation signal for 6 seconds at each monochromator step to ensure sufficient statistics in the line wings. To generate the spectrum, X-rays detected within a given region of interest centered on the expected energies of 3C and 3D were summed and projected onto the monochromator energy axis. These scans were repeated 20 times, and the resulting (summed) data is shown in Fig. 1. Each individual scan is fitted with a Voigt profile, a convolution of Gaussian and Lorentzian distributions. The Gaussian component arises from the Doppler width of the trapped ions and the apparatus profile of the monochromator (Hoesch et al., 2022). Meanwhile, the Lorentzian component, as explained in detail in our previous work (Kühn et al., 2022; Shah et al., 2024), results not only from the natural linewidth of the transitions but also from a pseudo-Lorentzian instrumental component caused by X-ray diffraction at beamline components Follath and Balzer (2010). Total intensities were determined from the area under the curve, derived through a maximum-likelihood fit of Voigt profiles using the Cash statistic Cash (1979); Kaastra (2017). Furthermore, 3C and 3D intensities were corrected for the presence of a 500-nm Al filter in front of SDDs, and normalized by the photon flux measured downstream of the EBIT with a calibrated photodiode, which together increase the ratio by 0.5%. Given that the intensity of the 3C and 3D transitions excited by the monochromatic X-ray beam is directly proportional to the oscillator strength of each transition Bernitt et al. (2012); Träbert (2022), we derived an oscillator strength ratio of 2.51(2) from the measured intensities. We note that the inner-shell satellite C of the Na-like ion, previously a major source of systematic uncertainties in many experiments on Fe XVII, does not affect the Ni XIX 3C/3D line ratio. This is primarily because line C is clearly separated from line 3D, falling well outside the scan range we used. From our low-resolution measurement with a 1-mm exit slit, we determined the difference eV.
The measured line ratio could be affected by periodic fluctuations of the actual photon energy around its nominal value, stemming from incorrect interpolation tables for the grating and mirror angular encoders (Follath and Balzer, 2010; Krempaský et al., 2011) of the monochromator, as discussed in our previous work Kühn et al. (2022); Shah et al. (2024); Togawa et al. (2023). To avoid this issue in subsequent works, we simultaneously measured a proxy for the fluctuations of the true photon energy with a photoelectron-energy spectrometer (ASPHERE) (Rossnagel et al., 2001). Unfortunately, this instrument was not available during the present measurements, but in follow-up campaigns Shah et al. (2024) we found with it oscillating differences between nominal and true photon energy of up to in the Ni XIX 3C and 3D scan range. We simulated such monchromator fluctuations with mock 3C and 3D Voigt profiles, resulting in systematic uncertainties of approximately 3% in the 3C/3D intensity ratio. A slight line profile asymmetry caused by X-ray diffraction at beamline components was quantified by fitting with skewed Voigt profiles, showing changes in the ratio of less than 0.5%. We observed scan-to-scan variations in the amplitudes of 3C and 3D of 3% which are not attributable to any known cause, and which we therefore take as a term in our systematic error budget. All above systematic uncertainties, including the 0.1% uncertainty in the 3C/3D ratio arising from a 10% uncertainty in the thickness of the 500-nm Al optical filter, have been taken into account in the final error budget for the inferred oscillator-strength ratio, namely , as displayed in Fig. 2. Please note that the circular polarization of the photon beam does not influence the ratio, as both transitions 3C and 3D have identical angular emission characteristics (see Appendix A in Ref. Steinbrügge et al. (2022))
To determine the natural linewidths of 3C and 3D, we used the Gaussian and Lorentzian widths extracted through Voigt fits to twenty scans of 3C and 3D. While Gaussian widths of approximately 50 meV for 3C and 3D were consistent with the experimental conditions, the extracted Lorentzian widths of 30 and 17 meV for 3C and 3D transitions, respectively, were clearly larger than expected from theory. We attribute this discrepancy to an additional pseudo-Lorentzian component induced by X-ray diffraction at beamline components Follath and Balzer (2010); Krempaský et al. (2011) that artificially raises the apparent Loretzian linewidth of observed transitions, as shown in our previous work Kühn et al. (2022). We checked this here by measuring the K, K, and K X-ray transitions from helium-like F VII and Ne IX several times. Their theoretical natural linewidths were taken from the NIST Atomic Spectral Database (ASD) Kramida et al. (2022); Cann and Thakkar (1992); Johnson et al. (2002), and we assigned them a conservative 10% uncertainty Gu et al. (2004). As Lorentzian contributions add linearly, in contrast to the quadratic addition of Gaussian widths, we subtracted the theoretical natural linewidths from the Lorentzian linewidths inferred from the measurements. We noticed an increase in the pseudo-Lorentzian beamline component dependent on the monochromator energy, which we model empirically as a quadratic function, determining the beamline Lorentzian contribution at 3D and 3C line energies to be 10.0(1.0) and 10.3(1.1) meV, respectively. We also derived beamline Lorentzian contributions of 7.2(5) meV and 7.4(5) meV at Fe XVII 3D and 3C energies, respectively, and upon comparison with the 7.0(3) meV obtained in our earlier work Kühn et al. (2022), which relied exclusively on the single F VIII K line, we find reasonably good agreement between the present and Kühn et al. (2022).
To derive the natural linewidths of Ni XIX 3C and 3D, we subtracted the pseudo-Lorentzian instrumental component from the Lorentzian widths obtained from Voigt fits to each scan of both lines. The determined natural linewidths were 19.8(1.2) and 7.1(1.0) meV for 3C and 3D, respectively, indicated as ”Method 1” in Tab. 1 and in Fig. 3. Their uncertainties include the statistical error on individual widths obtained from the fit, and systematic uncertainties from asymmetric line shapes (2%), monochromator energy fluctuations at 3C (2.2%) and 3D (2.3%) and at the helium-like K-shell reference transitions of helium-like F and Ne ions (6%).
Natural linewidths () | Line 3C (meV) | Line 3D (meV) | |
Method 1 | 19.8 1.2 | 7.1 1.0 | 0.10 |
Method 2 | 21.2 2.7 | 8.3 1.2 | 0.94 |
Mean (equal weights) | 20.5 1.7 | 7.7 0.7 | 0.63 |
Alternatively, we can use the following equations to determine the individual natural linewidths of 3C and 3D using the difference in their uncalibrated Lorentzian widths, as shown in Kühn et al. (2022):
However, in this second method, we have assumed that the pseudo-Lorentzian contribution from the beamline is similar for both 3C and 3D lines, as in Kühn et al. (2020). This allows the subtraction of Lorentzian widths obtained directly from the 3C and 3D Voigt fits ( meV). By utilizing this difference along with the measured (3C/3D) oscillator-strength ratio in the present work and the transition-energy (3C/3D) ratio measured in Gu et al. (2007) in the above equations, we determined the natural linewidths of 3C and 3D to be 21.2(2.7) and 8.3(1.2) meV, respectively. The final errors in this method (Method 2) account for Voigt fit statistical errors on individual widths, asymmetric line shapes (2%), effects of monochromator energy fluctuations on 3C (2.2%) and 3D (2.3%), and on (5%), in addition to the uncertainty in from this work and the uncertainty from Gu et al. (2007).
In both methods, we propagated uncertainties and covariances using Monte-Carlo methods, obtaining final natural linewidths values for 3C and 3D lines as unweighted means of all 20 individual measurements. A comparison of the results of both methods is displayed in Fig. 3(a). It shows that results for both linewidths are largely uncorrelated in Method 1, as it is less prone to systematics, while Method 2 exhibits a Pearson correlation coefficient close to one for both linewidths, indicating high correlation and greater susceptibility to systematics arising from amplitude ratios, energy ratios, and the assumption of similar beamline components. As an unbiased compromise, we choose their unweighted mean values as our final results presented in Tab. 1 and shown as magenta ellipses in Fig. 3(b).
III Results and Discussions
Line 3C | Line 3D | |||
Experiment | Theory | Experiment | Theory | |
20.50 1.70 | 20.867 0.026 | 7.70 0.70 | 7.570 0.027 | |
2.17 0.18 | 2.204 0.003 | 0.84 0.08 | 0.828 0.003 | |
3.12 0.26 | 3.170 0.004 | 1.17 0.11 | 1.150 0.004 | |
32.11 2.68 | 31.54 0.040 | 85.48 7.84 | 86.95 0.310 |
Figure 2 compares our results with earlier experimental data, observations, and predictions. Our own calculations using large-scale configuration interaction (large-CI) Cheung et al. (2021) and FAC-CI Gu (2008) methods, both accounting for the relativistic Breit interaction and quantum electrodynamics (QED) effects, yield an oscillator-strength ratio which agrees with our experiment within the uncertainties. Details of large-CI computations are given in Appendix A. Additional comparisons with other theoretical values from the literature showed typical departures within a 1–2 range, with the largest one from Zhang and Sampson (1989).
Three laboratory measurements of the 3C/3D intensity ratio of Ni XIX based on electron-impact excitation (EIE) were reported Brown et al. (2001c); Gu et al. (2004); Chen et al. (2006). The collisional excitation ratios from all three previous experiments span 1.90–2.35, slightly lower than the oscillator strength ratio measured in our experiment, but all four experiments are mutually consistent within uncertainties. Observations of the solar corona Hutcheon et al. (1976); McKenzie et al. (1980) and Capella Behar et al. (2001) show some scatter in the 3C/3D ratio, but in the absence of uncertainty estimates we cannot quantify the level of agreement with laboratory measurements.
Figure 3 (a) shows a comparison between measured 3C and 3D natural linewidths obtained using two different methods, which agree (see dashed diagonal lines) with each other and our predictions. Both our calculations using large-CI and FAC-CI methods agree within the 1 uncertainty of our experimental results. As can be seen in Figure 3 (b), the other older calculations from Zhang and Sampson (1989) and Safronova et al. (2001) show disagreement, however, their values just fall outside of the 2-sigma ellipse of mean natural linewidths, whereas somewhat recent calculations presented in Refs. Dong et al. (2003); Jönsson et al. (2014); Santana et al. (2015) show agreement within the 1-sigma ellipse.
The individual oscillator strengths of 3C and 3D can be derived from the natural linewidth using the relation , where . and represent the statistical weights of the initial () and final states (), respectively. The is the transition wavelength given in nanometers taken from Gu et al. (2007), and is the reduced Planck constant. Moreover, the experimental lifetime and transition rates can also be derived from the natural linewidth measurements. All of these measured quantities agree very well with our predictions, given in Tab. 2.
Our present experimental validation of oscillator strengths effectively eliminates uncertainties in atomic data as potential contributors to the observed discrepancies in nickel opacity measurements Nagayama et al. (2019) and iron opacity Bailey et al. (2007); Fontes et al. (2015); Kühn et al. (2022). The measured oscillator strengths offer direct application in astrophysical modeling, enabling the diagnosis of turbulent velocity in moderately optically thick plasmas Xu et al. (2002b); Werner et al. (2009). Upon comparing with oscillator strength ratios and natural linewidths from established databases such as NIST-ASD Kramida et al. (2022), AtomDB Foster et al. (2012), Chianti Del Zanna and Ishikawa (2009), and SPEX Kaastra et al. (1996), significant discrepancies were identified with our experimental results. There is a pressing need for updating these databases with such precise measurements in view of the crucial role of accurate data in modeling observational spectra from existing missions like Chandra and XMM-Newton, as well as upcoming observations with XRISM Tashiro et al. (2018), which has recently been successfully launched, as well as future planned missions such as Athena (Pajot et al., 2018), LEM (Kraft et al., 2022), HUBS (Cui et al., 2020), Arcus (Heilmann et al., 2022), HiReX (Nicastro et al., 2021), and Lynx (Schwartz et al., 2019). The present combination of experimental benchmark and converged calculations is a crucial consistency check for atomic data required to interpret X-ray observations in astrophysics, fusion, and high-energy density plasma research.
Acknowledgements.
This research was funded by the Max Planck Society (MPG) and the German Federal Ministry of Education and Research (BMBF) under project 05K13SJ2. C.S. acknowledges support from MPG and NASA under grant number 80GSFC21M0002. F.S.P. and M.A.L. acknowledge support from the NASA Astrophysics Program. The theoretical work has been supported by the US NSF Grant No. PHY-2012068, No. PHY-2309254, US Office of Naval Research Grant No. N00014-20-1-2513, and by the European Research Council (ERC) under the Horizon 2020 Research and Innovation Programme of the European Union (Grant Agreement No. 856415). Calculations were performed in part using the computing resources at the University of Delaware, in particular the Caviness and DARWIN high-performance computing clusters. We acknowledge DESY (Hamburg, Germany), a member of the Helmholtz Association HGF, for the provision of experimental facilities. Parts of this research were carried out at PETRA III. We thank Jens Viefhaus and Rolf Follath for valuable discussions on X-ray monochromator resolution and performance, and the synchrotron-operation team and P04 team at PETRA III for their skillful and reliable work. MSS thanks MPIK, Heidelberg for hospitality.*
Appendix A Theory
The calculations are carried out using a large-scale configuration interaction (CI) method Cheung et al. (2021), including correlations from all ten electrons. Basis sets of increasing sizes are used to check for convergence of the values. The basis set is designated by the highest principal quantum number for each partial wave included. For example, [] means that all orbitals up to are included for the partial waves. We begin by considering all possible single and double excitations to orbitals up to from the , even and , , odd configurations, correlating 8 electrons. We verified that inclusion of the even and , odd configurations as basic configurations has a negligible effect on the energies and relevant matrix elements. The calculated contributions to the energies of Ni XIX are listed in Tab. 3. The results are compared with a revised analysis of the experimental data Kramida (2019). We use -coupling and NIST-style LS-coupling term designations for comparisons. Contributions to the reduced matrix elements and and the oscillator strengths ratios are listed in Table 4. The energy ratio is 1.018 and the oscillator strength ratio is 2.64(2).
To assess the impact of triple excitations, we consider a broad range of configurations up to . As demonstrated in Tables 3 and 4, these excitations result in negligible corrections to both energies and matrix elements. Subsequently, we expand the basis set to , leading to a significant improvement in the agreement of energies with experimental values and a minimal shift in the ratio (-0.008). Further accounting of contributions from the shell with the basis improves agreement with experimental energies, albeit with a marginal contribution to the ratio (-0.006). A comparison of results for and obtained in length and velocity gauges reveals only a marginal difference of 0.001 for the basis. Expanding the basis set to produces a modest improvement in energies compared to experiment, accompanied by a slight shift in the ratio (-0.001). Further expansion to results in minor corrections to energies, with an even smaller contribution to the ratio (-0.0005). The quantum electrodynamic (QED) contributions are included following Ref. Tupitsyn et al. (2016). The QED inclusion has small effect on the individual line energies, however, a considerable contribution to the energy difference of 3C and 3D, see Tab. 3. Furthermore, the ratio is changed by -0.01 by accounting for QED.
Additionally, we compute the transition rates for all other transitions contributing to the radiative decay of the 3 and 3 levels. The sums of these rates are small and listed in Tab. 4. Linewidth values correspond to total transition rates, with uncertainties computed from the largest uncertainties in the 3C and 3D matrix elements, including additional configurations in CI space and QED. Additional uncertainties due to orbitals for transition rates and linewidths, based on recent studies of Fe16+ Shah et al. (2024), are included, and the final uncertainties are computed based on the relative difference.
Configuration | Expt. Kramida (2019) | Triples | QED | Final | Diff. Kramida (2019) | Diff. Kramida (2019) | |||||||
0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | 0 | |||
1 | 7381990 | 7374679 | -206 | 3579 | 351 | 797 | 372 | 44 | 7379615 | 2375 | 0.03% | ||
2 | 7409915 | 7403138 | -2 | 2880 | 264 | 724 | 331 | 29 | 7407364 | 2551 | 0.03% | ||
3 | 7431735 | 7424692 | -4 | 3023 | 261 | 738 | 340 | 102 | 7429152 | 2583 | 0.03% | ||
1 | 7440050 | 7433248 | -11 | 2836 | 277 | 730 | 332 | 85 | 7437497 | 2553 | 0.03% | ||
2 | 7105260 | 7096413 | 17 | 3417 | 478 | 751 | 350 | 1052 | 7102477 | 2783 | 0.04% | ||
1 | 7122600 | 7114019 | 15 | 3303 | 415 | 722 | 337 | 1052 | 7119862 | 2738 | 0.04% | ||
1 | 7247700 | 7249597 | 14 | 3386 | 498 | 735 | 345 | 1403 | 7255978 | -8278 | 0.11% | ||
1 | 7807700 | 7801245 | 19 | 2336 | 409 | 684 | 363 | 72 | 7805128 | 2572 | 0.03% | ||
2 | 7825770 | 7819486 | 19 | 1839 | 415 | 651 | 359 | 87 | 7822856 | 2914 | 0.04% | ||
4 | 7825280 | 7819623 | 19 | 2044 | 410 | 677 | 362 | 88 | 7823223 | 2057 | 0.03% | ||
3 | 7830930 | 7825368 | 17 | 1687 | 414 | 633 | 353 | 81 | 7828554 | 2376 | 0.03% | ||
2 | 7847100 | 7841657 | 18 | 1638 | 410 | 654 | 356 | 86 | 7844819 | 2281 | 0.03% | ||
3 | 7857640 | 7852407 | 17 | 1705 | 407 | 630 | 350 | 90 | 7855606 | 2034 | 0.04% | ||
1 | 7901400 | 7899252 | 3 | 1681 | 384 | 638 | 352 | 136 | 7902446 | -1046 | 0.01% | ||
2 | 7972475 | 7967013 | 17 | 1683 | 484 | 657 | 362 | 432 | 7970647 | 1828 | 0.02% | ||
2 | 7980810 | 7975017 | 17 | 1933 | 475 | 667 | 362 | 427 | 7978897 | 1913 | 0.02% | ||
3 | 7986640 | 7981013 | 16 | 1759 | 481 | 638 | 356 | 443 | 7984706 | 1934 | 0.02% | ||
1 | 8041800 | 8040754 | -29 | 1668 | 383 | 633 | 353 | 372 | 8044132 | -2332 | 0.03% | ||
(eV) | 17.4074 | 17.5440 | -0.0039 | -0.0016 | -0.0002 | -0.0006 | -0.0001 | 0.0291 | 17.5669 |
0.30012 | 0.18530 | 2.670 | |||||
Triples | 0.29999 | -0.00013 | 0.18530 | 0.00000 | 2.668 | -0.002 | |
0.30031 | 0.00019 | 0.18568 | 0.00038 | 2.662 | -0.008 | ||
0.30060 | 0.18582 | 2.663 | |||||
+ | 0.30018 | -0.00013 | 0.18581 | 0.00013 | 2.656 | -0.006 | |
0.30032 | 0.00001 | 0.18571 | 0.00003 | 2.662 | -0.001 | ||
0.30031 | -0.00001 | 0.18572 | 0.00001 | 2.661 | -0.0005 | ||
QED | -0.00012 | 0.00028 | -0.010 | ||||
Final | 0.29993 | 0.18613 | 2.64(2) | ||||
Recommended transition rate (s-1) | 3.168(4) | 1.148(4) | |||||
Other transitions (s-1) | 1.75 | 1.59 | |||||
Total rate (s-1) | 3.170(4) | 1.150(4) | |||||
Linewidth (meV) | 20.867(26) | 7.570(27) |